### v1

Abstract
Given two points in the plane, and a set of “obstacles” given as curves through the plane with assigned weights, we consider the point-separation problem, which asks for the minimum-weight subset of the obstacles separating the two points. A few computational models for this problem have been previously studied. We give a unified approach to this problem in all models via a reduction to a particular shortest-path problem, and obtain improved running times in essentially all cases. In addition, we also give fine-grained lower bounds for many cases.
keywords:
obstacle separation, point separation, geometric intersection graph, $Z_{2}$-homology, fine-grained lower bounds
category:
[…]
1 Introduction
Figure 1: An instance of the $(s,t)$ point-separation problem.
Given points $s$ and $t$ in the plane, and a weighted set of obstacles $\mathcal{C}$ defined by simple closed curves (possibly also including their interiors), the $(s,t)$ point-separation problem asks for the minimum-weight subset $C$ of $\mathcal{C}$ such that any path from $s$ to $t$ intersects some obstacle in $C$. Equivalently, $s$ and $t$ are in different connected components of $\mathbb{R}^{2}\setminus(\cup_{\gamma\in C}\gamma)$. We say such a subset separates $s$ and $t$. An example of this problem can be found in fig. 1.
This is a natural problem that arises in various scenarios. As a toy application of this problem: Suppose every night your dog runs from his doghouse to your backpack to eat your homework, taking an unpredictable route (making use of windows and doors). You have noticed that your dog will forget about your homework if it smells a treat. You have a number of candidate locations to place treats, and you’d like to ensure your dog is distracted from your homework every day using the fewest treats possible. See fig. 2 for this example. Similar applications also arise when considering security (e.g., replacing the batteries in the fewest number of your dead security cameras to cover all possible paths from the entrance to your bank vault). Additionally, $(s,t)$ point-separation has found an application as a tool for constant-factor approximation of the well-studied APX-hard problem “barrier-resilience” [KumarLSS21].
LawnWindowWindowWindowWindowWindowWindowWindowWindowWindowWindowWindowWindowWindowWindowKitchenLiving RoomBathroomBedroomCouchDoghouseBackpackLawnWindowWindowWindowWindowWindowWindowWindowWindowWindowWindowWindowWindowWindowWindowKitchenLiving RoomBathroomBedroomCouchDoghouseBackpack
Item Eating locations ( ) Walls Couch Windows
Cost 1 0 0 $\infty$
Figure 2: An example of the point-separation problem applied to placing distracting doggy treats around a house, with the minimum-weight solution on the right.
For brevity, we will henceforth say “curve” to mean a simple closed curve in $\mathbb{R}^{2}\setminus\{s,t\}$. The algorithmic complexity of the $(s,t)$ point-separation problem depends on the chosen computational model of the obstacles. Previous works use a few different models. We categorize and name the different classes of models used as follows:
  • Specific Obstacle-Type Models: Assume the set of curves $\mathcal{C}$ takes on a special form with a standard representation, such as a set of disks, circles, or line segments.
  • The Oracle-based Intersection Graph Model: Assume the existence of several $O(1)$-time oracles that would allow the computation of the intersection graph $G$ of $\mathcal{C}$ (the graph over vertex set $\mathcal{C}$ whose edges correspond to pairwise intersections), as well as some additional information for each edge related to $s$ and $t$ (detailed in section 2).
  • The Arrangement Model: Assume the arrangement of $\mathcal{C}$ is provided as input, in the form of a plane (multi-)graph, with the faces corresponding to $s$ and $t$ labelled. This is a more graph-theoretic formulation. See fig. 3 for an example of such an arrangement.
Figure 3: An arrangement representation of the instance in fig. 1.
Prior works discussing this problem each focused on only one of these paradigms, and the names for each paradigm are our own. We will not assume general position, so no paradigm is strictly more general than the others, since there are arrangements of $n$ obstacles with $\Theta(n^{2})$ pairwise intersections, but only $O(n)$ unique intersection points.
In our work, our key method is to frame all models of this problem in terms of a “homology cover”. Homology is a very broad topic in algebraic topology, which we will not attempt to summarize in this paper. We aim our paper at a typical computational geometry audience, and we will not assume prior knowledge of homology. We will present the necessary aspects in section 2.
1.1 Prior Work (Brief)
In this subsection, we very briefly outline some key aspects of prior work. A significantly extended form of this section is in LABEL:sec:prior-work.
Most importantly for our methods, Kumar, Lokshtanov, Saurabh and Suri [KumarLSS21, Section 6] describe an algorithm that runs in polynomial time in the arrangement model. Their algorithm implicitly makes use of the homology cover (perhaps unintentionally), in the same sense that we will use it. In fact, their algorithm is some ways analogous to an algorithm of Chambers, Erickson, Fox, and Nayyeri [ChambersEFN23] for minimum-cuts (and maximum-flow) in surface graphs. These two algorithms are the main inspiration for our approach to the $(s,t)$ point-separation at a high-level, in all models.
1.2 Results and Organization
We obtain improved algorithms for the $(s,t)$-point separation problem in several cases, which we outline in table 1. As mentioned earlier, all of our positive results make use of a reduction to a shortest-path problem in the so-called “homology cover”. We discuss this formulation in section 2, and then again more rigorously in LABEL:sec:homology-formal. Using this reduction, the upper bounds are then given in section 3.
Model Weights Old New Cond. L.B.
Oracle Yes $O(n^{3})$ [CabelloG16] $\widetilde{O}(n^{(3+\omega)/2})$ (Thm. 3.3) $\Omega^{*}\left(n^{2}\right)$ (Thm. LABEL:thm:lb-segments-clique)
Oracle No $O(n^{3})$ [CabelloG16] $O(n^{\omega}\log n)$ (Thm. 3.5) $\Omega^{*}(n^{3/2})$ (Cor. LABEL:thm:unweighted-lb)
Arrangement Yes $O(km^{2}\lg k)$ [KumarLSS21] $O(km+k^{2}\lg k)$ (Thm. 3.7) $\Omega^{*}\left(km\right)^{\dagger}$ (Cor. LABEL:cor:lb-polylines)
Segments No $O(n^{3})$ [CabelloG16] $O(n^{7/3}\log^{1/3}n)$ (Cor. 3.12) $\Omega^{*}(n^{3/2})$ (Thm. LABEL:thm:unweighted-lb)
Axis-aligned
segments
No $O(n^{3})$ [CabelloG16] $O(n^{2}\log\log n)$ (Cor. 3.12) None
Unit Disks No $O(n^{2}\log^{3}n)$ [CabelloM18] $O(n^{2}\log n)$ (Cor. 3.12) None
Disks No $O(n^{3})$ [CabelloG16] $O(n^{2}\log n)$ (Cor. 3.12) None
$O(1)$-length
polylines
No $O(n^{3})$ [CabelloG16] $O(n^{7/3}\log^{1/3}n)$ (Cor. 3.12) $\Omega^{*}\left(n^{3/2}\right)$ (Thm. LABEL:thm:unweighted-lb)
$O(1)$-length
rectilinear
polylines
No $O(n^{3})$ [CabelloG16] $O(n^{2}\log\log n)$ (Cor. 3.12) $\Omega^{*}(n^{3/2})$ (Thm. LABEL:thm:unweighted-lb)
Segments Yes $O(n^{3})$ [CabelloG16] $\widetilde{O}(n^{7/3})$ (Cor. LABEL:cor:separation-biclique-cover-times) $\Omega^{*}\left(n^{2}\right)$ (Thm. LABEL:thm:lb-segments-clique)
Axis-aligned
segments
Yes $O(n^{3})$ [CabelloG16] $\widetilde{O}(n^{2})$ (Cor. LABEL:cor:separation-biclique-cover-times) None
$O(1)$-length
polylines
Yes $O(n^{3})$ [CabelloG16] $\widetilde{O}(n^{7/3})$ (Cor. LABEL:cor:separation-biclique-cover-times) $\Omega^{*}\left(n^{2}\right)$ (Thm. LABEL:thm:lb-segments-clique)
$O(1)$-length
rectilinear
polylines
Yes $O(n^{3})$ [CabelloG16] $\widetilde{O}(n^{2})$ (Cor. LABEL:cor:separation-biclique-cover-times) $\Omega^{*}(n^{2})$ (Thm. LABEL:thm:lb-segments-clique)
Table 1: The time complexities of our algorithms for various obstacle models. In all cases, $n:=|\mathcal{C}|$. For the arrangement model, $k$ denotes the vertex count of the arrangement, and $m$ is the number of obstacle-vertex incidences (so $m\geq k$). The notation $\Omega^{*}(\cdot)$ hides sub-polynomial factors. The notation denotes that this is only true for one particular mutual dependence of $k$, $m$, and $n$ (and is a slight abuse of notation). $\omega$ is the matrix-multiplication exponent ($\omega<2.371339$ [alman2024asymmetryyieldsfastermatrix]).
We also obtain several fine-grained lower bounds in LABEL:sec:lower-bounds. Our lower bounds also all have a shared foundation, which will be stated in LABEL:thm:k-cycle-reduction. The resulting lower bounds are based on a few different hypotheses, and we give their details in LABEL:sec:lb-details.
2 Homology and Obstacles [Informal]
In this section, we will explain the main topological tools we will use to approach the $(s,t)$ point-separation problem. However, this will be an information section, not requiring prior knowledge of any aspect of topology. Rather, we will give an equivalent formulation of the important pieces using simpler tools from computational geometry. We give a more formal treatment in LABEL:sec:homology-formal Throughout this section, we will make (implicit) assumptions of general position and such, but LABEL:sec:homology-formal does not need these.
At a high-level, there are two main steps to the constructions of our approach. First, we will discuss what it means for a simple curve to separate $s$ and $t$, and what tools exist to classify such curves. Second, we will discuss what it means for a set of obstacles to separate $s$ and $t$. That is, when the union of the obstacles contains a simple curve separating $s$ and $t$¿
The simplest possible case of asking whether a simple curve separates $s$ and $t$ is characterized by the point-in-polygon problem, which asks: Given a point $p$ and a (simple) polygon $P$, is $p$ inside $P$? In fact, this is equivalent to asking if $P$ separates $p$ and the point at infinity. There is a folklore algorithm for this problem that takes any ray $r$ starting at $p$, and counts the number of intersections between $r$ and $P$. Then, $P$ contains $p$ if and only if the number of intersections is odd. In fact, essentially the same algorithm can be used to test whether or not $P$ separates two points $p_{1}$ and $p_{2}$. Rather than using a ray, take the segment $\overline{p_{1}p_{2}}$. Count the number of intersections between $\overline{p_{1}p_{2}}$ and $P$. The count is odd if and only if $P$ separates $p_{1}$ and $p_{2}$. Moreover, there is in fact nothing special about rays or line segments in this problem. Any (closed) path between $p_{1}$ and $p_{2}$ would cross $P$ the same number of times, modulo $2$. See fig. 4 for examples for these algorithms.
Figure 4: A demonstration of the algorithm for the point-in-polygon problem (left), as well as the problem of testing whether or not a polygon separates two points (middle). On the right, alternative paths for the separation problem are given.
All three of these algorithms are equivalent in a sense. In fact, there is a further generalization: Given two points $s$ and $t$ on the sphere, a simple and closed curve $C$ (not covering $s$ or $t$), and any $s-t$ path $\pi$, $C$ separates $s$ and $t$ if and only if $\pi$ crosses $C$ an odd number of times (regardless of the choice of $\pi$). Since the extended plane is homeomorphic to the sphere, a ray in the plane corresponds to a (simple) path in the sphere. See fig. 5 for an example.
Figure 5: Examples of point-pairs separated or not separated by a given closed path between the pair, and the corresponding intersections of those curves.
The problems we have discussed so far are all completely static, and the methods do not provide much structure for solving more difficult problems. One of the most common ways to extend the point-in-polygon problem is to fix the polygon (or curve) $P$, and aim to support fast queries of points. This problem is known as “point-location”, and it is well-studied in computational geometry [BergCKO08] However, we want a different sort of structure: We have a fixed pair of points $s$ and $t$, and we wish to classify the curves that separate them. Since we have fixed $s$ and $t$, we can also fix the path $\pi$ between them – in most cases, we will use the line segment $\overline{st}$. Then, the problem of classifying curves that separate $s$ and $t$ becomes the problem of classifying curves that cross $\overline{st}$ an odd number of times.
It will be helpful for demonstration to make some transformations to the space we work in. That is, we will perform a sequence of homeomorphisms. We start with the extended plane with the two marked points $s$ and $t$, and the line segment $\overline{st}$ (see fig. 9). No curves we will be classifying cross $s$ or $t$, so we may assume there are punctures at $s$ and $t$. We can expand these punctures with a homeomorphism (see fig. 9). Next, since we have the extended plane, we can make one of the punctures the outer face, obtaining an annulus (see fig. 9). In performing these steps, the line segment $\overline{st}$ becomes a path between the inner and outer boundaries of the annulus.
{subfigure}
[t]0.28 Refer to caption {subfigure}[t]0.28 Refer to caption {subfigure}[t]0.28 Refer to caption
Figure 6: $s$ and $t$ are punctures.
Figure 7: $s$ and $t$ are large punctures.
Figure 8: $s$ is the outer face, and the space is an annulus.
Figure 9: A demonstration of how the (extended) plane with two punctures is homeomorphic to the annulus.
We now present a method for constructing an important space called the homology cover111 We use the term “homology cover” to refer to the one-dimensional $\mathbb{Z}_{2}$-homology cover of the annulus. . Take the specified path between the two boundaries (see fig. 13) and slice it open (see fig. 13). Then, create a second copy of the sliced annulus, and glue them together in a way that matches up the orientations of the sliced ends (see fig. 13).
{subfigure}
[t]0.28 Refer to caption {subfigure}[t]0.28 Refer to caption {subfigure}[t]0.28 Refer to caption
Figure 10: An annulus with a specified path.
Figure 11: The annulus is sliced open at the specified path.
Figure 12: Two copies of the sliced annulus are glued together.
Figure 13: The “cut and glue” construction of the homology cover, as well as how it maps a closed curve that separates the two boundaries (blue) and one that does not (red).
Alternatively, an equivalent construction is to create two copies of the original space (whether that be the extended plane or the annulus), and use each side of the path from $s$ to $t$ as a (separate) “portal” between the two copies. A more general form of this “portal” idea has been studied in the form of “portalgons” [loffler_et_al, ophelders2024shortest], of which the homology cover is essentially a special case. See fig. 14 for an example of this construction.
Figure 14: The “portal” construction of the homology cover. Each colour (or dot/dash pattern) is a single closed curve in the homology cover. The two solid lines are the portals.
One important aspect of this construction is that it involves the connection of two identical copies of the annulus (i.e., the extended plane with punctures $s$ and $t$). In fig. 13, we also show how this construction transforms two curves – one separating the two boundaries, and one not separating them. The structure we will make use of to study curves separating $s$ and $t$ in the plane ultimately stems from an important set of facts:
[…]
For a simple curve $C$ in the annulus (or the plane) that gets mapped to the set $C^{\prime}$ in the homology cover, and a point $p$ along $C$ that gets mapped to corresponding points $p_{1}$ and $p_{2}$ in the homology cover, the following are all equivalent:
  • $C$ separates $s$ and $t$.
  • $C^{\prime}$ separates the two boundaries in the homology cover.
  • $C^{\prime}$ has one connected component (i.e., it is one closed curve instead of two).
  • The points $p_{1}$ and $p_{2}$ are connected by a path through $C^{\prime}$.
It is this last characterization that will be the most critical for obstacles. In particular, a consequence of this characterization is that if two corresponding points $p_{1}$ and $p_{2}$ in the homology cover (corresponding in the sense that they are identical points in different copies of the annulus/plane) have a simple path $\pi$ between them, then that path can be mapped to a closed curve separating $s$ and $t$.
2.1 Obstacles and the Homology Cover
We now have tools for characterizing closed curves that separate $s$ and $t$. We’d now like to characterize sets of obstacles that separate $s$ and $t$. In other words, we’d like to characterize unions of closed curves (obstacles) that contain a closed curve separating $s$ and $t$. We’ll give two different tools for this, first for the arrangement model, then for the oracle model. For simplicity, we will work exclusively with obstacles that do not individually separate $s$ and $t$ (that is, obstacles that get mapped to two separate closed curves in the homology cover). We will reduce to this case algorithmically at the start of section 3.
In the arrangement model of the $(s,t)$ point-separation problem, we are given a plane graph $D$ (the arrangement) with specified faces $s$ and $t$ (can be obtained from points via point-location), and a set of obstacles $\sigma$, each given as a connected subgraph. The homology cover has a simple graph-theoretic interpretation in this framework: Take any (simple) dual-path $\pi$ of faces from $s$ to $t$ – this is will serve as the “portal”. Next, create two copies $D_{1},D_{2}$ of $D$, each with the faces $s$ and $t$ removed. We will create a combined graph $\overline{D}$ starting from the union of $D_{1}$ and $D_{2}$. For each edge $e=uv$ used in the dual-path $\pi$, delete $e$ from both $D_{1}$ and $D_{2}$ (inside $\overline{D}$). Denote the corresponding vertices of $u$ and $v$ in each of $D_{1}$ and $D_{2}$ as $u_{1},v_{1}$ and $u_{2},v_{2}$, respectively. Then create new edges $u_{1}v_{2}$ and $u_{2}v_{1}$ in $\overline{D}$. This form of the homology cover is visualized in fig. 15.
Figure 15: The homology cover of the arrangement given in fig. 3. The purple edges are the modified “crossing” edges.
With this particular arrangement structure Kumar, Lokshtanov, Saurabh and Suri [KumarLSS21, Section 6] built a graph using one vertex per obstacle-vertex incidence (in each copy of the plane graph). We will build a smaller graph of similar form to theirs. First, for each obstacle $C\in\sigma$ (which induces a subgraph of $D$), pick some arbitrary canonical point $x\in C$. Since this choice is arbitrary, it will sometimes be useful to assume it is a specific point – this will primarily be useful for specific obstacle types. Assume that $x$ is also a vertex in the subgraph of $D$ induced by $C$ (and if it is not, modify $D$ so that it is, with either a bisection or a new degree-$1$ vertex). Note that since $C$ is connected in $D$, every other vertex $x^{\prime}\in C$ is connected to $x$ by only edges in $C$. Given a plane graph $D$ with faces $s$ and $t$ a dual-path $\pi$, obstacles $\sigma$, and canonical points for each obstacle, the auxiliary graph $H$ is a bipartite graph constructed as follows:
  • The first set of vertices are the copies of arrangement vertices in the homology cover. For a vertex $v\in V(D)$, we denote these two copies $v^{+}$ and $v^{-}$.
  • The second set of vertices are the copies of the obstacles/canonical points in the homology cover. For an obstacle $\gamma\in\sigma$, we denote these two copies as $(\gamma,-)$ and $(\gamma,+)$.
  • A vertex copy $v^{b_{1}}$ has an edge connecting it to a canonical point copy $(\gamma,b_{2})$ when:
    • $v\in\gamma$ (that is, the point $v$ is covered by the obstacle $\gamma$ in the arrangement), and…
    • The canonical point $x$ of $\gamma$ has a path through the edges of $\gamma$ to $v$ whose intersection with the edges of $\pi$ is size $b_{1}+b_{2}$ modulo $2$. Equivalently, the projected connected component of $\gamma$ containing the canonical point copy $(\gamma,b_{2})$ also contains $v^{b_{1}}$.
The auxiliary graph is useful because the shortest-path between any two corresponding vertex copies or (separately) any two corresponding canonical point copies both correspond exactly to the solution to the $(s,t)$ point-separation problem. The high-level construction to prove this fact is as follows:
  • Suppose there is a set of obstacles $C$. We’d like to determine when $C$ separates $s$ and $t$.
  • Denote the set of copies of the obstacles in $C$ in the homology cover as $C^{\prime}$, so that every element of $C^{\prime}$ is a closed curve representing one of the two connected components of an element of $C$ projected into the homology cover.
  • By section 2, we deduce that $C$ separates $s$ and $t$ if and only if there is a path from $v^{+}$ to $v^{-}$ along the union of obstacle copies in $C^{\prime}$, for some arrangement vertex $v$ among pairs of elements in $C$ only.
  • Moreover, since each obstacle is connected, we may further assume that any such path visits each obstacle copy at most once, and moreover that it only visits each obstacle at most once.
  • We may further assume that any such path visits the canonical point copy of each such obstacle copy, by inserting a path to and then from the canonical point copy while the obstacle copy is visited. Note that the resulting path may not be simple in the plane.
  • Such paths also directly correspond to paths in the auxiliary graph, using only “obstacle vertices” corresponding to elements of $C$ and the “arrangement vertices” incident to pairs of elements in $C$.
  • With appropriate weights, the problem of finding the minimum-weight set $C$ with this property then reduces to the problem of finding the pair $v^{+},v^{-}$ with the shortest distance in the auxiliary graph, so this is a shortest-path problem!
  • A similar argument can show that an alternative algorithmic formulation is to find the pair $(\gamma,-),(\gamma,+)$ with the shortest-distance in the auxiliary graph.
This essentially completes the set of tools necessary for the arrangement model. These arguments are given in more detail in LABEL:sec:homology-formal.
Figure 16: The intersection graph (left) and the intersection graph in the homology cover (right) for the curves given in fig. 1. The purple edges are the “crossing” edges.
For the oracle model, the purpose of the oracles will be to construct the geometric intersection graph of the obstacle copies in the homology cover (which we will denote as $\overline{G}$). We will call $\overline{G}$ the intersection graph in the homology cover. See fig. 16 for an example. We will present two different constructions of this graph. They are equivalent, and each of them is quite simple, but they will serve two different purposes: The first will clarify which types of oracle queries are necessary, and provide an algorithmic construction. The second will instead build on the tools we have for the arrangement model, proving the correctness of another shortest-paths approach.
Let $\mathcal{C}$ denote the set of obstacles in the plane. We assume the canonical points for each obstacle are given (or implied) – they will be used by the oracle. We also assume some simple $s-t$ path $\pi$ is fixed – this will also be used by the oracle. As mentioned In most cases, $\pi$ will be $\overline{st}$. The first construction of the graph is as follows:
  • For each obstacle $\gamma\in\mathcal{C}$, create two vertices $(\gamma,-)$ and $(\gamma,+)$.
  • For each pair of obstacles $\gamma_{1},\gamma_{2}$ with canonical points $x_{1},x_{2}$, and values $b_{1},b_{2}\in\{-,+\}$, we connect vertices $(\gamma_{1},b_{1})$ and $(\gamma_{2},b_{2})$ by an edge if there is a path from $x_{1}$ to $x_{2}$ in $\gamma_{1}\cup\gamma_{2}$ crossing $\pi$ exactly $k$ times, for some $k$ where $k\equiv b_{1}+b_{2}\pmod{2}$.
Testing for this condition is actually the only type of oracle query needed (although we will also require support for the case when $\gamma_{1}=\gamma_{2}$ later). Cabello and Giannopoulos [CabelloG16] used a larger set of query types, but together they can be used to perform this type by carefully choosing the canonical points and fixing $\pi$ to $\overline{st}$, so our variant of the oracle model slightly more general (although practically equivalent).
We’d like a similar algorithmic property for the intersection graph in the homology cover $\overline{G}$ that we have for the auxiliary graph $H$. The second construction will be what gives us that property, by constructing $\overline{G}$ from $H$:
  • For each arrangement vertex $v^{b}$ in $H$, let its adjacent vertices be denoted $x_{1},\dots,x_{k}$. Delete $v^{b}$ and create a clique over $\{x_{1},\dots,x_{k}\}$.
  • After performing this for every arrangement vertex $v^{b}$, the result is exactly the intersection graph in the homology cover $\overline{G}$.
  • Therefore, the set of vertices visited by the shortest path from some $(\gamma,-)$ to $(\gamma,+)$ (over all possible $\gamma\in\mathcal{C}$) corresponds to the minimum-weight set of obstacles separating $s$ and $t$.
For specific obstacle types, we will usually work with $\overline{G}$, although we will do it implicitly. Hence, with the tools from this section, we can now discuss algorithms in all three model types as solutions to a certain form of shortest-path problem.
3 Algorithmic Results
In this section, we will devise algorithms for the $(s,t)$ point-separation problem based on our homology cover structures. We will devise algorithms for all three model types: Some that work with the arrangements of curves, some that work with the “oracle model” (the intersection graph in the homology cover), and some that work with specific types of obstacles. Most of our algorithms are fairly simple reductions to various known shortest-path algorithms, greatly simplifying some of the previous approaches to the $(s,t)$ point-separation problem. However, some of them are more involved.
In the previous section, we assumed no obstacle individually separated $s$ and $t$. Before moving on, we quickly show this is enough: {observation} Let $\mathcal{C}$ be a set of obstacles, and let $s$ and $t$ be points. Suppose an oracle exists that, for an obstacle $\gamma\in\mathcal{C}$, determines whether $\gamma$ itself separates $s$ and $t$, all in $O(1)$ time. Let $\mathcal{C}_{0}$ be the set of all obstacles that do. Then an instance of the weighted (unweighted) $(s,t)$ point-separation problem over $\mathcal{C}$ can be reduced to an instance of the weighted (unweighted) $(s,t)$ point-separation problem over $\mathcal{C}\setminus\mathcal{C}_{0}$ in $O(|\mathcal{C}|)$ time. In particular, such a reduction returns the best solution out of the solved subproblem (if any), and each of the individual obstacles forming solutions (if any).
3.1 General Weighted Obstacles in Oracle Model
With no significant further work, we already obtain some naïve algorithms by using LABEL:lemma:sp (or the constructions in section 2):
Theorem 3.1.
For points $s$ and $t$, and a weighted set of obstacles $\mathcal{C}$ that have $k$ pairwise intersections, the $(s,t)$ point-separation problem can be solved in $O(|\mathcal{C}|\cdot k+|\mathcal{C}|^{2}\log|\mathcal{C}|)$ time.
Proof 3.2.
Compute the intersection graph in the homology cover $\overline{G}$ of $\mathcal{C}$. For each $\gamma\in\mathcal{C}$, compute the shortest-path from $(\gamma,-)$ to $(\gamma,+)$ in $\overline{G}$. The smallest such path induces the solution by LABEL:lemma:sp. Each shortest-path can be computed in $O(k+|\mathcal{C}|\log|\mathcal{C}|)$ time by Dijkstra’s algorithm with a Fibonacci heap [clrs], and there are $O(|\mathcal{C}|)$ total such paths.
Note that theorem 3.1 matches the result of [CabelloG16], although the algorithm is much simpler. When $k\in\Theta(|\mathcal{C}|^{2})$ (i.e., dense intersection graphs), this algorithm runs in $O(|\mathcal{C}|^{3})$ time, which could also be obtained by running Floyd-Warshall for APSP instead of Dijkstra. It is not known that the general form of weighted APSP can be solved in “truly” subcubic time (see LABEL:subsec:apsp). However, since the weights are applied to vertices, not edges, we can obtain a faster algorithm via the results from the discussion in LABEL:subsec:apsp:
Theorem 3.3.
For points $s$ and $t$, and a weighted set of obstacles $\mathcal{C}$, the $(s,t)$ point-separation problem can be solved in $\widetilde{O}(|\mathcal{C}|^{(3+\omega)/2})$ time in the oracle model, where $2.3713392\leq\omega<$ 2.371339 is the matrix multiplication exponent.
Proof 3.4.
Apply the same reduction to APSP as before, but use the algorithm of Abboud, Fischer, Jin, Williams, and Xi [abboud2025allpairs] for APSP with real vertex weights.
3.2 General Unweighted Obstacles in Oracle Model
We are also interested in the $(s,t)$ point-separation problem for unit weights, which corresponds to shortest-paths for unit weights. In this case, we can obtain faster algorithms:
Theorem 3.5.
For points $s$ and $t$, and an unweighted set of obstacles $\mathcal{C}$ that have $k$ pairwise intersections, the $(s,t)$ point-separation problem can be solved in $O(|\mathcal{C}|^{\omega}\log|\mathcal{C}|)$ time, where $2.3713392\leq\omega<$ 2.371339 is the matrix multiplication exponent.
Proof 3.6.
Compute the intersection graph in some homology cover $\overline{G}$ of $\mathcal{C}$, and then run unweighted undirected APSP [Seidel95] on $\overline{G}$, for $O(|\mathcal{C}|^{\omega}\log|\mathcal{C}|)$ total time.
3.3 General Weighted Obstacles in Arrangement Model
We can also obtain results using the arrangement instead of the intersection graph. The simplest of these is a slight modification of theorem 3.1:
Theorem 3.7.
For points $s$ and $t$, a weighted set of obstacles $\mathcal{C}$, and an arrangement $D,\sigma$ of $\mathcal{C}$, where $\sigma$ is given as lists of vertices, let $m=\sum_{c\in\sigma}|c|$ be the total number of obstacle-vertex incidences. Then, the $(s,t)$ point-separation problem can be solved in $O(\min(|\mathcal{C}|,|V(D)|)\cdot m+\min(|\mathcal{C}|,|V(D)|)^{2}\log\min(|\mathcal{C}|,|V(D)|))$ time.
Proof 3.8.
Let $\overline{H}$ be the auxiliary graph in the homology cover, which has $O(m)$ vertices and edges. By LABEL:lemma:sp, it suffices to do one of the following:
  • Compute the SSSP tree from each obstacle in $\overline{H}$.
  • Compute the SSSP tree from each arrangement vertex in $\overline{H}$.
Let $k=\min(|\mathcal{C}|,|V(D)|)$. Each SSSP computation can be performed in $O(m+k\log k)$ time, and $k$ such computations suffices to solve the problem.
This theorem in particular is of note because as we will see later, it is essentially optimal assuming the APSP conjecture, at least in the case when $|\mathcal{C}|=\Theta(|V(D)|^{2})$ and $m=\Theta(|\mathcal{C}|)$, as we will see in LABEL:sec:lower-bounds.
3.4 Restricted Obstacle Classes without Weights
As we will discuss in LABEL:subsec:apsp, Chan and Skrepetos [Chan07, Chan10] studied APSP for several forms of unweighted/undirected geometric intersection graphs. Their intersection graphs are in the plane, but with some extra work it is possible to study intersection graphs in the homology cover, and consequently the unweighted $(s,t)$ point-separation problem.
Theorem 3.9.
[…]
Proof 3.10.
Run APSP for the intersection graph in the homology cover via LABEL:thm:unweighted-intersection-apsp.
To use this result, we need to algorithms for static intersection in the homology cover:
Lemma 3.11.
[…]
General Disks $O(n\log n)$
Axis-Aligned Line Segments $O(n\log\log n)$
Arbitrary Line Segments $O(n^{4/3}\log^{1/3}n)$
The proof of the lemma is left to LABEL:sec:static-intersection-hc. At a high-level, all cases are first reduced to the planar static intersection problem. In the case of line segments, this becomes the standard planar static intersection problem for line segments. For disks, this is not the case, and the algorithm is more involved.
The combination of these two results give an important corollary:
Corollary 3.12.
For points $s$ and $t$, the unweighted $(s,t)$ point-separation problem can be solved in $O(n^{2}\log\log n)$ time for axis-aligned line segments (or $O(1)$-length rectilinear polylines), $O(n^{7/3}\log^{1/3}n)$ time for line segments (or $O(1)$-length polylines), and $O(n^{2}\log n)$ time for general disks or circles that do not contain both $s$ and $t$.
3.5 Restricted Obstacle Classes with Weights
We will now present a method for solving the weighted $(s,t)$ point-separation problem using a tool called “biclique covers”, which are essentially a tool for a type of graph sparsification.
For a graph $G=(V,E)$, a
Definition 1.
biclique in $G$ is a complete bipartite subgraph ($A\times B\subset E$, where $A,B\subset V$ are disjoint). The

### v2

@@
 Abstract
-Given two points in the plane, and a set of “obstacles” given as curves through the plane with assigned weights, we consider the point-separation problem, which asks for the minimum-weight subset of the obstacles separating the two points.
+Given two points in the plane, and a set of “obstacles” given as curves through the plane with assigned weights, we consider the point-separation problem, which asks for a minimum-weight subset of the obstacles separating the two points.
 A few computational models for this problem have been previously studied.
@@
 category:
-[…]
 1 Introduction
 Figure 1: An instance of the $(s,t)$ point-separation problem.
-Given points $s$ and $t$ in the plane, and a weighted set of obstacles $\mathcal{C}$ defined by simple closed curves (possibly also including their interiors), the $(s,t)$ point-separation problem asks for the minimum-weight subset $C$ of $\mathcal{C}$ such that any path from $s$ to $t$ intersects some obstacle in $C$.
+Given points $s$ and $t$ in the plane, and a weighted set of obstacles $\mathcal{C}$ defined by simple closed curves (possibly also including their interiors), the $(s,t)$ point-separation problem asks for a minimum-weight subset $C$ of $\mathcal{C}$ such that any path from $s$ to $t$ intersects some obstacle in $C$.
 Equivalently, $s$ and $t$ are in different connected components of $\mathbb{R}^{2}\setminus(\cup_{\gamma\in C}\gamma)$.
 We say such a subset separates $s$ and $t$.
-An example of this problem can be found in fig.
-1.
+An example of this problem can be found in Figure 1.
 This is a natural problem that arises in various scenarios.
@@
 You have a number of candidate locations to place treats, and you’d like to ensure your dog is distracted from your homework every day using the fewest treats possible.
-See fig.
-2 for this example.
+See Figure 2 for this example.
 Similar applications also arise when considering security (e.g., replacing the batteries in the fewest number of your dead security cameras to cover all possible paths from the entrance to your bank vault).
@@
   • Specific Obstacle-Type Models: Assume the set of curves $\mathcal{C}$ takes on a special form with a standard representation, such as a set of disks, circles, or line segments.
-  • The Oracle-based Intersection Graph Model: Assume the existence of several $O(1)$-time oracles that would allow the computation of the intersection graph $G$ of $\mathcal{C}$ (the graph over vertex set $\mathcal{C}$ whose edges correspond to pairwise intersections), as well as some additional information for each edge related to $s$ and $t$ (detailed in section 2).
+  • The Oracle-based Intersection Graph Model: Assume the existence of several $O(1)$-time oracles that would allow the computation of the intersection graph $G$ of $\mathcal{C}$ (the graph whose vertices are exactly the curves $\mathcal{C}$ and whose edges correspond to pairwise intersections of these curves), as well as some additional information for each edge related to $s$ and $t$ (detailed in Section 2).
   • The Arrangement Model: Assume the arrangement of $\mathcal{C}$ is provided as input, in the form of a plane (multi-)graph, with the faces corresponding to $s$ and $t$ labelled.
 This is a more graph-theoretic formulation.
-See fig.
-3 for an example of such an arrangement.
-Figure 3: An arrangement representation of the instance in fig. 1.
+See Figure 3 for an example of such an arrangement.
+Figure 3: An arrangement representation of the instance in Figure 1, as well as an arrangement of a smaller instance.
 Prior works discussing this problem each focused on only one of these paradigms, and the names for each paradigm are our own.
@@
 Their algorithm implicitly makes use of the homology cover (perhaps unintentionally), in the same sense that we will use it.
-In fact, their algorithm is some ways analogous to an algorithm of Chambers, Erickson, Fox, and Nayyeri [ChambersEFN23] for minimum-cuts (and maximum-flow) in surface graphs.
+In fact, their algorithm is in some ways analogous to an algorithm of Chambers, Erickson, Fox, and Nayyeri [ChambersEFN23] for minimum-cuts (and maximum-flow) in surface graphs.
 These two algorithms are the main inspiration for our approach to the $(s,t)$ point-separation at a high-level, in all models.
@@
 Rather, we will give an equivalent formulation of the important pieces using simpler tools from computational geometry.
-We give a more formal treatment in LABEL:sec:homology-formal Throughout this section, we will make (implicit) assumptions of general position and such, but LABEL:sec:homology-formal does not need these.
+We give a more formal treatment in LABEL:sec:homology-formal.
+Throughout this section, we will make (implicit) assumptions of general position and such, but the deferred formal treatment does not need these.
 At a high-level, there are two main steps to the constructions of our approach.
@@
 Second, we will discuss what it means for a set of obstacles to separate $s$ and $t$.
-That is, when the union of the obstacles contains a simple curve separating $s$ and $t$¿
+That is, when does the union of the obstacles contain a simple curve separating $s$ and $t$?
 The simplest possible case of asking whether a simple curve separates $s$ and $t$ is characterized by the point-in-polygon problem, which asks: Given a point $p$ and a (simple) polygon $P$, is $p$ inside $P$?
@@
 Any (closed) path between $p_{1}$ and $p_{2}$ would cross $P$ the same number of times, modulo $2$.
-See fig.
-4 for examples for these algorithms.
+See Figure 4 for examples for these algorithms.
 Figure 4: A demonstration of the algorithm for the point-in-polygon problem (left), as well as the problem of testing whether or not a polygon separates two points (middle).
@@
 Since the extended plane is homeomorphic to the sphere, a ray in the plane corresponds to a (simple) path in the sphere.
-See fig.
-5 for an example.
+See Figure 5 for an example.
 Figure 5: Examples of point-pairs separated or not separated by a given closed path between the pair, and the corresponding intersections of those curves.
@@
 One of the most common ways to extend the point-in-polygon problem is to fix the polygon (or curve) $P$, and aim to support fast queries of points.
-This problem is known as “point-location”, and it is well-studied in computational geometry [BergCKO08] However, we want a different sort of structure: We have a fixed pair of points $s$ and $t$, and we wish to classify the curves that separate them.
+This problem is known as “point-location”, and it is well-studied in computational geometry [BergCKO08].
+However, we want a different sort of structure: We have a fixed pair of points $s$ and $t$, and we wish to classify the curves that separate them.
 Since we have fixed $s$ and $t$, we can also fix the path $\pi$ between them – in most cases, we will use the line segment $\overline{st}$.
@@
 That is, we will perform a sequence of homeomorphisms.
-We start with the extended plane with the two marked points $s$ and $t$, and the line segment $\overline{st}$ (see fig.
-9).
+We start with the extended plane with the two marked points $s$ and $t$, and the line segment $\overline{st}$ (see Figure 6(a)).
 No curves we will be classifying cross $s$ or $t$, so we may assume there are punctures at $s$ and $t$.
-We can expand these punctures with a homeomorphism (see fig.
-9).
-Next, since we have the extended plane, we can make one of the punctures the outer face, obtaining an annulus (see fig.
-9).
+We can enlarge these punctures into holes with a homeomorphism.
+Next, since we have the extended plane, we can make one of the punctures the outer face, obtaining an annulus (see Figure 6(b) and Figure 6(c)).
 In performing these steps, the line segment $\overline{st}$ becomes a path between the inner and outer boundaries of the annulus.
-{subfigure}
-[t]0.28 Refer to caption {subfigure}[t]0.28 Refer to caption {subfigure}[t]0.28 Refer to caption
-Figure 6: $s$ and $t$ are punctures.
-Figure 7: $s$ and $t$ are large punctures.
-Figure 8: $s$ is the outer face, and the space is an annulus.
+(a) $s$ and $t$ are punctures, $\pi$ is a path between them.
+(b) $s$ and $t$ are enlarged to become holes, and the space is an annulus.
+(c) The two sides of $\pi$ can be coloured to aid later notation.
 Figure 6: A demonstration of how the (extended) plane with two punctures is homeomorphic to the annulus.
@@
 .
-Take the specified path between the two boundaries (see fig.
-13) and slice it open (see fig.
-13).
-Then, create a second copy of the sliced annulus, and glue them together in a way that matches up the orientations of the sliced ends (see fig.
-13).
-{subfigure}
-[t]0.28 Refer to caption {subfigure}[t]0.28 Refer to caption {subfigure}[t]0.28 Refer to caption
-Figure 10: An annulus with a specified path.
-Figure 11: The annulus is sliced open at the specified path.
-Figure 12: Two copies of the sliced annulus are glued together.
+Take the specified path $\pi$ between the two boundaries (see Figure 7(a)) and slice it open (see Figure 7(b)).
+Then, create a second copy of the sliced annulus, and glue them together in a way that matches up the orientations of the sliced ends (see Figure 7(c)).
+(a) An annulus with a specified path $\pi$.
+(b) The annulus is sliced open along the specified path $\pi$.
+(c) Two copies of the sliced annulus are glued together.
 Figure 7: The “cut and glue” construction of the homology cover, as well as how it maps a closed curve that separates the two boundaries (blue) and one that does not (red).
@@
 A more general form of this “portal” idea has been studied in the form of “portalgons” [loffler_et_al, ophelders2024shortest], of which the homology cover is essentially a special case.
-See fig.
-14 for an example of this construction.
+See Figure 8 for an example of this construction.
 Figure 8: The “portal” construction of the homology cover. Each colour (or dot/dash pattern) is a single closed curve in the homology cover. The two solid lines are the portals.
 One important aspect of this construction is that it involves the connection of two identical copies of the annulus (i.e., the extended plane with punctures $s$ and $t$).
-In fig.
-13, we also show how this construction transforms two curves – one separating the two boundaries, and one not separating them.
+In Figure 7, we also show how this construction transforms two curves – one separating the two boundaries, and one not separating them.
 The structure we will make use of to study curves separating $s$ and $t$ in the plane ultimately stems from an important set of facts:
@@
 Then create new edges $u_{1}v_{2}$ and $u_{2}v_{1}$ in $\overline{D}$.
-This form of the homology cover is visualized in fig.
-15.
-Figure 15: The homology cover of the arrangement given in fig. 3. The purple edges are the modified “crossing” edges.
+This form of the homology cover is visualized in Figure 9.
+(a) The arrangement with a specified path $\pi$.
+(b) A simple separating cycle in the arrangement.
+(c) The edges of the arrangement crossing $\pi$ are deleted.
+(d) Two copies of are created, and the deleted edges are added as “crossing” edges (in purple).
+(e) A path between two copies of the same original vertex that would project onto the simple separating cycle in the original space from Figure 9(b).
+Figure 9: The construction of the homology cover for the second arrangement given in Figure 3.
 With this particular arrangement structure Kumar, Lokshtanov, Saurabh and Suri [KumarLSS21, Section 6] built a graph using one vertex per obstacle-vertex incidence (in each copy of the plane graph).
@@
 Since this choice is arbitrary, it will sometimes be useful to assume it is a specific point – this will primarily be useful for specific obstacle types.
-Assume that $x$ is also a vertex in the subgraph of $D$ induced by $C$ (and if it is not, modify $D$ so that it is, with either a bisection or a new degree-$1$ vertex).
+Assume that $x$ is also a vertex in the subgraph of $D$ induced by $C$ (and if it is not, modify $D$ so that it is, with either a subdivision or a new degree-$1$ vertex).
 Note that since $C$ is connected in $D$, every other vertex $x^{\prime}\in C$ is connected to $x$ by only edges in $C$.
-Given a plane graph $D$ with faces $s$ and $t$ a dual-path $\pi$, obstacles $\sigma$, and canonical points for each obstacle, the auxiliary graph $H$ is a bipartite graph constructed as follows:
+Given a plane graph $D$ with faces $s$ and $t$, a dual-path $\pi$, obstacles $\sigma$, and canonical points for each obstacle, the auxiliary graph $H$ is a bipartite graph constructed as follows:
   • The first set of vertices are the copies of arrangement vertices in the homology cover. For a vertex $v\in V(D)$, we denote these two copies $v^{+}$ and $v^{-}$.
   • The second set of vertices are the copies of the obstacles/canonical points in the homology cover. For an obstacle $\gamma\in\sigma$, we denote these two copies as $(\gamma,-)$ and $(\gamma,+)$.
+  • The values -- and $+$ in each of the above are indicator bits, and can be stored as 00 and $1$, respectively.
   • A vertex copy $v^{b_{1}}$ has an edge connecting it to a canonical point copy $(\gamma,b_{2})$ when:
-    • $v\in\gamma$ (that is, the point $v$ is covered by the obstacle $\gamma$ in the arrangement), and…
-    • The canonical point $x$ of $\gamma$ has a path through the edges of $\gamma$ to $v$ whose intersection with the edges of $\pi$ is size $b_{1}+b_{2}$ modulo $2$.
+    • $v\in\gamma$ (that is, the point $v$ belongs to the obstacle $\gamma$ in the arrangement), and…
+    • The canonical point $x$ of $\gamma$ has a path through the edges of $\gamma$ to $v$ whose intersection with the edges of $\pi$ has size $b_{1}+b_{2}$ modulo $2$ (i.e., if $b_{1}=b_{2}$, the size must be even, and if $b_{1}\neq b_{2}$, the size must be odd).
 Equivalently, the projected connected component of $\gamma$ containing the canonical point copy $(\gamma,b_{2})$ also contains $v^{b_{1}}$.
-The auxiliary graph is useful because the shortest-path between any two corresponding vertex copies or (separately) any two corresponding canonical point copies both correspond exactly to the solution to the $(s,t)$ point-separation problem.
+We visualize an example of the auxiliary graph in Figure 10.
+(a) A canonical point is added along each obstacle, and edges are created from the canonical point to each arrangement vertex along the obstacle, subject to the parity conditions.
+(b) A shortest-path between the two copies of an obstacle’s canonical point. The obstacles whose canonical points lie in this path form a set of obstacles separating $s$ and $t$.
+(c) The set of obstacles induced by the shortest path in Figure 10(b).
+Figure 10: The construction of the auxiliary graph for the homology cover given in Figure 9.
+The auxiliary graph is useful because the shortest path between any two corresponding vertex copies or (separately) any two corresponding canonical point copies both correspond exactly to the solution to the $(s,t)$ point-separation problem.
 The high-level construction to prove this fact is as follows:
@@
 This essentially completes the set of tools necessary for the arrangement model. These arguments are given in more detail in LABEL:sec:homology-formal.
-Figure 16: The intersection graph (left) and the intersection graph in the homology cover (right) for the curves given in fig. 1. The purple edges are the “crossing” edges.
+(a) The intersection graph with the path $\pi$.
+(b) The intersection graph after deleting the edges whose underlying paths cross $\pi$ an odd number of times.
+(c) Two copies of the intersection graph are connected with “crossing” edges (in purple).
+(d) A shortest-path in the homology cover between two copies of an obstacle.
+Figure 11: How the intersection graph can be transformed into the intersection graph in the homology cover.
 For the oracle model, the purpose of the oracles will be to construct the geometric intersection graph of the obstacle copies in the homology cover (which we will denote as $\overline{G}$).
 We will call $\overline{G}$ the intersection graph in the homology cover.
-See fig.
-16 for an example.
+See Figure 11 for an example.
 We will present two different constructions of this graph.
@@
 We also assume some simple $s-t$ path $\pi$ is fixed – this will also be used by the oracle.
-As mentioned In most cases, $\pi$ will be $\overline{st}$.
+In most cases, $\pi$ will be $\overline{st}$.
 The first construction of the graph is as follows:
@@
 Testing for this condition is actually the only type of oracle query needed (although we will also require support for the case when $\gamma_{1}=\gamma_{2}$ later).
-Cabello and Giannopoulos [CabelloG16] used a larger set of query types, but together they can be used to perform this type by carefully choosing the canonical points and fixing $\pi$ to $\overline{st}$, so our variant of the oracle model slightly more general (although practically equivalent).
+Cabello and Giannopoulos [CabelloG16] used a larger set of query types, but together they can be used to perform this type by carefully choosing the canonical points and fixing $\pi$ to $\overline{st}$, so our variant of the oracle model is slightly more general (although practically equivalent).
 We’d like a similar algorithmic property for the intersection graph in the homology cover $\overline{G}$ that we have for the auxiliary graph $H$.
@@
   • After performing this for every arrangement vertex $v^{b}$, the result is exactly the intersection graph in the homology cover $\overline{G}$.
-  • Therefore, the set of vertices visited by the shortest path from some $(\gamma,-)$ to $(\gamma,+)$ (over all possible $\gamma\in\mathcal{C}$) corresponds to the minimum-weight set of obstacles separating $s$ and $t$.
+  • Therefore, the set of vertices visited by the shortest path from some $(\gamma,-)$ to $(\gamma,+)$ (over all possible $\gamma\in\mathcal{C}$, breaking ties arbitrarily) corresponds to a minimum-weight set of obstacles separating $s$ and $t$.
 For specific obstacle types, we will usually work with $\overline{G}$, although we will do it implicitly.
@@
 Compute the intersection graph in the homology cover $\overline{G}$ of $\mathcal{C}$.
-For each $\gamma\in\mathcal{C}$, compute the shortest-path from $(\gamma,-)$ to $(\gamma,+)$ in $\overline{G}$.
+For each $\gamma\in\mathcal{C}$, compute the shortest path from $(\gamma,-)$ to $(\gamma,+)$ in $\overline{G}$.
 The smallest such path induces the solution by LABEL:lemma:sp.
-Each shortest-path can be computed in $O(k+|\mathcal{C}|\log|\mathcal{C}|)$ time by Dijkstra’s algorithm with a Fibonacci heap [clrs], and there are $O(|\mathcal{C}|)$ total such paths.
+Each shortest path can be computed in $O(k+|\mathcal{C}|\log|\mathcal{C}|)$ time by Dijkstra’s algorithm with a Fibonacci heap [clrs], and there are $O(|\mathcal{C}|)$ total such paths.
 Note that Theorem 3.1 matches the result of [CabelloG16], although the algorithm is much simpler.
 When $k\in\Theta(|\mathcal{C}|^{2})$ (i.e., dense intersection graphs), this algorithm runs in $O(|\mathcal{C}|^{3})$ time, which could also be obtained by running Floyd-Warshall for APSP instead of Dijkstra.
-It is not known that the general form of weighted APSP can be solved in “truly” subcubic time (see LABEL:subsec:apsp).
-However, since the weights are applied to vertices, not edges, we can obtain a faster algorithm via the results from the discussion in LABEL:subsec:apsp:
+It is not known if the general form of weighted APSP can be solved in “truly” subcubic time (see LABEL:subsec:apsp).
+However, since the weights are applied to vertices, not edges, there is a known faster algorithm (also discussed in LABEL:subsec:apsp):
 Theorem 3.3.
@@
 3.2 General Unweighted Obstacles in Oracle Model
-We are also interested in the $(s,t)$ point-separation problem for unit weights, which corresponds to shortest-paths for unit weights. In this case, we can obtain faster algorithms:
+We are also interested in the $(s,t)$ point-separation problem for unit weights, which corresponds to shortest paths for unit weights. In this case, we can obtain faster algorithms:
 Theorem 3.5.
-For points $s$ and $t$, and an unweighted set of obstacles $\mathcal{C}$ that have $k$ pairwise intersections, the $(s,t)$ point-separation problem can be solved in $O(|\mathcal{C}|^{\omega}\log|\mathcal{C}|)$ time, where $2.3713392\leq\omega<$ 2.371339 is the matrix multiplication exponent.
+For points $s$ and $t$, and an unweighted set of obstacles $\mathcal{C}$, the $(s,t)$ point-separation problem can be solved in $O(|\mathcal{C}|^{\omega}\log|\mathcal{C}|)$ time, where $2.3713392\leq\omega<$ 2.371339 is the matrix multiplication exponent.
 Proof 3.6.
@@
 Let $\overline{H}$ be the auxiliary graph in the homology cover, which has $O(m)$ vertices and edges. By LABEL:lemma:sp, it suffices to do one of the following:
-  • Compute the SSSP tree from each obstacle in $\overline{H}$.
+  • Compute the single-source shortest-path (SSSP) tree from each obstacle in $\overline{H}$.
   • Compute the SSSP tree from each arrangement vertex in $\overline{H}$.
-Let $k=\min(|\mathcal{C}|,|V(D)|)$. Each SSSP computation can be performed in $O(m+k\log k)$ time, and $k$ such computations suffices to solve the problem.
+Let $k=\min(|\mathcal{C}|,|V(D)|)$. Each SSSP computation can be performed in $O(m+k\log k)$ time, and $k$ such computations suffice to solve the problem.
 This theorem in particular is of note because as we will see later, it is essentially optimal assuming the APSP conjecture, at least in the case when $|\mathcal{C}|=\Theta(|V(D)|^{2})$ and $m=\Theta(|\mathcal{C}|)$, as we will see in LABEL:sec:lower-bounds.
@@
 Theorem 3.9.
-[…]
+For points $s$ and $t$, and an unweighted set of obstacles $\mathcal{C}$.
+Let $\overline{\mathcal{C}}$ be the set of $2|\mathcal{C}|$ “mapped” obstacles in the homology cover.
+Let $\text{SI}(n,m)$ (“static intersection”) be the time complexity for checking if each of $n$ different objects in $\overline{\mathcal{C}}$ intersects any object in some subset $C\subset\overline{\mathcal{C}}$ of size $m=|C|$.
+Assume $\text{SI}(n,m)$ is super-additive, so that $\text{SI}(n_{1},m_{1})+\text{SI}(n_{2},m_{2})\leq\text{SI}(n_{1}+n_{2},m_{1}+m_{2})$.
+Then $(s,t)$ point-separation over $\mathcal{C}$ can be solved in $O(n\text{SI}(n,n))$ time.
 Proof 3.10.
-Run APSP for the intersection graph in the homology cover via LABEL:thm:unweighted-intersection-apsp.
-To use this result, we need to algorithms for static intersection in the homology cover:
+Run APSP for the intersection graph in the homology cover via the algorithm of Chan [ChanS19, Theorem 2] (see further discussion in LABEL:subsec:apsp).
+To use this result, we need efficient algorithms for static intersection in the homology cover:
 Lemma 3.11.
-[…]
-[…]
+For points $s$ and $t$, the following values of $\text{SI}(n,n)$ hold for restricted obstacle types in the homology cover:
+Obstacle Class $\text{SI}(n,n)$
 General Disks $O(n\log n)$